OpenStudy (anonymous):

During a recent rainstorm the basement of the police station was flooded. A pump was started, but after 2 hours it was decided that the process was taking too long. At that point a second pump was hooked up, and the job was completed using both pumps in another 3 hours. Working separately, it would have taken the original pump 4 hours longer than the new one. How long would it have taken the original pump?

OpenStudy (legends):

Ok i THINK ITS 5.5 HOURS.

OpenStudy (legends):

3/2=1.5+4=5.5

OpenStudy (anonymous):

Thanks!

OpenStudy (legends):

No problem^